OpenStudy (anonymous):

Solve the system of equations: 3x + y = -8 -2x - y = 6

OpenStudy (matt101):

You could just add the two equations: 3x + y = -8 + -2x - y = 6 ----------------- x = -2 Then plug x into either of the original equations to find y. For example: 3(-2) + y = -8 -6 + y = -8 y = -2 Therefore the solution is (-2,-2).

OpenStudy (anonymous):

thanks

hero (hero):

I just posted that solution already. The same exact thing